OpenStudy (phucdatbich69):

Johan found that the equation –2|8 – x| – 6 = –12 had two possible solutions: x = 5 and x = –11. Which explains whether his solutions are correct?

OpenStudy (phucdatbich69):

Will fan and medal!

OpenStudy (mrmudd184):

ok

OpenStudy (mrmudd184):

He is not correct because he made a sign error.

OpenStudy (rcp031):

–2|8 – x| – 6 = –12 -2|8 – x|=-6 |8 – x|=3 8-x=-3 & 8-x=3 solve for x